How do you determine whether the discontinuity is removable or non removable? Show with an example.
1
Expert's answer
2020-04-28T18:50:44-0400

Removal discontinuity and non removable discontinuity exist in a rational function.

Removable Discontinuity:

Suppose we have a rational function "f(x)" which is discontinuous at some point say "x = a" but if the limit "\lim_{x \to a } f(x)" exists then we can say that the function has a removable discontinuity at "x = a"

Example:

Let us assume that

"f(x) = \frac{x^2 - 9}{x+3}"

Observing the function, we can say that the function has discontinuity at "x = -3" but the limit of the function at "x \to -3"

is given by

"\lim_{x \to -3} \frac{x^2 - 9}{x + 3} \\
= \lim_{x \to -3} \frac{(x+3)(x-3)}{x+3} \hspace{1 cm}" [ "\because a^2 - b^2 = (a+b)(a-b)" ]

"= \lim_{x \to -3} (x-3) \hspace{1 cm}" [ By canceling common factors "(x+3)" from both numerator and denominator ]

"= -3 - 3 \hspace{1 cm}" [ By substituting "x = - 3" into the limit function ]

"= -6"

Hence, the limit of the function exists and "x = - 3" is the removable discontinuity of the function "f(x)"

"\\"

Non Removable Discontinuity:

In non removable discontinuity, after canceling the common factors from both numerator and denominator, the function still has a discontinuity at some point.

Example:

Let us assume that

"g(x) = \frac{x^2 - x - 2}{x^2 - 5x - 6}"

"\hspace{0.5 cm} \,\, \,\,= \frac{x^2 - 2x + x - 2}{x^2 - 6x + x - 6 }"

"\hspace{0.5 cm} \,\,\,\, = \frac{x(x-2) + 1(x-2)}{x(x-6) +1(x-6)}"

"\hspace{0.5 cm} \,\,\,\, = \frac{(x-2)(x+1)}{(x-6)(x+1)}"

"\hspace{0.5 cm} \,\,\,\, = \frac{x-2}{x-6}"

After canceling the common factors "(x+1)" from both numerator and denominator, we can see that the function "g(x)" still has discontinuity at "x = 6 ." This is non removable discontinuity.